OpenStudy (anonymous):

How do I solve t^2+t-28=0, I'm doing Completing the Square.

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

yeah, but I am having trouble factoring out the answers...

OpenStudy (anonymous):

@shatteredskittlez where r u getting stuck at ? would u like to show your work?

OpenStudy (anonymous):

Sure! t^2+t-28=0 +28 +28 t^2+t=28 Now I would square the 1 in 1t (bx in the equation) (1/2)^2 0.5^2 0.25 t^2+t+0.25=28+0.25 thta was as far as i got before i got confused on factoring my answer...

OpenStudy (anonymous):

@shatteredskittlez u r going right continuing from where u left (t+0.5)^2 = 28.25

OpenStudy (anonymous):

i used the identity (a+b)^2 = a^2 + 2*a*b+ b^2

OpenStudy (anonymous):

ah ok! thx for ur help! :D Problem solved!